OpenStudy (waheguru):

WHat is cscA = 1.1547 So far I know that 1/sinA = 1.1547 but how do I go about from here?

jimthompson5910 (jim_thompson5910):

1/sinA = 1.1547 1 = 1.1547*sinA ... Multiply both sides by sinA. 1/1.1547 = sinA ... Divide both sides by 1.1547 0.8660 = sinA sinA = 0.8660 A = ???

jimthompson5910 (jim_thompson5910):

or you could use the shortcut that if 1/x = y then x = 1/y basically you take the reciprocal of both sides

OpenStudy (waheguru):

Is it 60 Degrees?

jimthompson5910 (jim_thompson5910):

roughly, yes

OpenStudy (waheguru):

thankyou

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (waheguru):

@jim_thompson5910 Just want to make sure I understood this correctly so secA = 1.0403 A = 16 (rounded)

jimthompson5910 (jim_thompson5910):

yes correct